OpenStudy (anonymous):

write a recursive formula for the sequence 7 4 1 -2 -5

OpenStudy (anonymous):

And then find the next term.

OpenStudy (anonymous):

-8 is the next term. There is a common difference of 3.

OpenStudy (anonymous):

You could write the formula as: (current term) = (previous term) - 3

OpenStudy (anonymous):

or \[x_i = x_{i-1} - 3\] where i is your current index

OpenStudy (anonymous):

there's these choices: an=an-1+3, where a1=7; -8 an=an-1+3, where a1=-3; 22 an=an-1-3, where a1=7; -8 an=an-1-3, where a1= -8; 7

OpenStudy (anonymous):

Well the formula I wrote is the same as the formula in options C and D. \[a_1 = 7\] because that is the first term in your sequence

OpenStudy (anonymous):

So it's C.

OpenStudy (anonymous):

\[a_n = -8\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome